# Reduction to master integrals and transverse integration identities

### v1

@@ 1 Introduction
In this paper, we study the impact of transverse integration (TI) identities on the solution of the Laporta system of equations, and assess the performance improvement they can provide. Starting from an integral family contributing to a process, we identify integrals corresponding to diagrams with either fewer external legs than the original family, or that can be factorized into lower-loop integrals. These can be mapped to simpler integral families having fewer invariants or fewer loops, making the identities for these new families much simpler. We map these integrals to the new families via a transverse integration, namely a tensor decomposition in the subspace that is orthogonal to their external legs. While this technique already appeared, e.g. in the context of integrand decomposition Mastrolia:2016dhn and multi-loop numerical unitarity Ita:2015tya , its impact on a traditional IBP reduction is unexplored.
@@ 3 The main idea in a pedagogical example
Its kinematics is described by $3$ massless and one massive external momenta and we have $3$ independent external invariants, as follows
@@ 3.1 Integrals with fewer external legs
As we will see in section 4, the form factors $C_{j}$ can be obtained by contracting the tensor integral on the l.h.s. of (27) with projectors $P_{j,\perp}$ that are linear combinations of the same tensors $T_{j,\perp}$. The coefficients of this linear combinations are rational functions of $D_{\perp}$ that can be determined in a general way and only depend on the tensor rank $\beta_{1}+\beta_{2}$ and not on the integral family (see also Anastasiou:2023koq ). By applying these projectors to the tensor integrals, we obtain scalar products of the form $(k_{i,\perp}\cdot k_{j,\perp})$, that in turn may be rewritten as
@@ 4.1 Sectors with fewer external legs
where $C_{j}$ are scalar form factors and $T_{\perp,j}^{\mu_{1}\cdots\mu_{r}}$ is a basis of rank-$r$ covariant tensors in the $(D-D_{\perp})$-dimensional transverse space. Such a basis can be chosen using
@@
while the others can be obtained from $T_{1,\perp}$ via independent permutations of the Lorentz indexes. Here $g_{\perp}^{\mu\nu}$ is the projection of the metric tensor into the subspace that is transverse to $q_{1},\ldots,q_{D_{\parallel}}$, as defined above. The form factors $C_{j}$ are computed by contracting the l.h.s. of eq. (73) with suitable projectors $P_{j,\perp}$ such that
@@ 4.2 Factorizable sectors
$I_{F;a_{1},\cdots a_{t}0\cdots 0}=\left(\int\prod_{j=1}^{\ell_{1}}d^{D}k_{j}\frac{1}{D_{1}^{a_{1}}\cdots D_{t_{1}}^{a_{t_{1}}}}\right)\left(\int\prod_{j=\ell_{1}+1}^{\ell_{1}+\ell_{2}}d^{D}k_{j}\frac{1}{D_{1}^{a_{t_{1}+1}}\cdots D_{t_{1}}^{a_{t_{1}+t_{2}}}}\right),$ (85)
@@
$I_{F;a_{1},\cdots a_{t}0\cdots 0}[N(k_{1},\ldots,k_{\ell})]=\left(\int\prod_{j=\ell_{1}+1}^{\ell_{1}+\ell_{2}}d^{D}k_{j}\frac{1}{D_{1}^{a_{t_{1}+1}}\cdots D_{t_{1}}^{a_{t_{1}+t_{2}}}}\right)I_{F_{1};a_{1}\cdots a_{t}0\cdots 0}[N(k_{1},\ldots,k_{\ell})],$ (87)
where the numerator $N$ is polynomial in the ISPs of $F$ and therefore in the scalar products one can build from the loop and external momenta of $F$. We thus focus on the $\ell_{1}$-loop integral $I_{F_{1};a_{1}\cdots a_{t}0\cdots 0}[N(k_{1},\ldots,k_{\ell})]$ and handle its numerator by following the same TI steps as in section 4.1. More specifically, $N(k_{1},\ldots,k_{\ell})$ may contain scalar products which cannot be written as a linear combination of generalized denominators of $F_{1}$. The corresponding integrals, however, will be cast – by means of the algorithm in 4.1 – as a linear combination of integrals $I_{F;\vec{b}}$ in the family $F_{1}$. Moreover, by reviewing these steps of the TI algorithm, it is easy to see that the coefficients $N_{\vec{b}}$ of such linear combination will contain, in their denominators, only scalar products between the external legs $q_{1},\ldots,q_{D_{\parallel}}$ of sector $S_{F_{1};a_{1}\cdots a_{t_{1}}0\cdots 0}$. Hence, the dependence of these coefficients on any other momentum which may appear in $N(k_{1},\ldots,k_{\ell})$, including the other loop momenta $k_{\ell_{1}+1},\ldots k_{\ell_{1}+\ell_{2}}$, will be in the numerator of these coefficients. We thus obtain a relation of the form
@@
The same method can be used, recursively, for factorizable sectors whose corner integral is the product of more than two one-loop factors.
@@ 5.2 Implementation
These three identities will thus be combined numerically over finite fields. While this strongly reduces the size of the expressions, it doesn’t completely fix the issue for complex multi-scale integrals having high rank. This happens in particular for step 1 above, and sometimes also for step 3. We note, however, that the identities in eq. (70) and eq. (79) for individual generalized denominators or scalar products are relatively simple. Moreover, they are simply substitutions of loop variables which we can implement at the integrand level. When these are raised to high powers or multiplied to become high-rank integrals, on the other hand, expressions can still become very large. Finite fields technologies, however, allow us to fix this issue with our second trick. We first encode the expressions for the required generalized denominators and scalar products as (linear) polynomials in the loop variables appearing on the r.h.s. of the equations we use in steps 1 and 3 above. The coefficients of such polynomials, which only depend on the external variables, are instead treated numerically over finite fields. This drastically simplifies their expressions. Hence, for each numerical evaluation, we can afford to build higher powers of (products of) these polynomials via a sequence of polynomial multiplications over finite fields. At the time of writing, FiniteFlow does not expose polynomial multiplication algorithms to its Mathematica interface, so these have been implemented as a sequence of linear substitutions, for which FiniteFlow’s sparse solver was used. A more specialized and optimized algorithm might be used in the future, but this was sufficient for our applications. These two tricks thus allowed us to implement complex TI identities without ever dealing with huge expressions.
@@ 5.3 Examples
In particular, we report on the relative performance improvement of a numerical evaluation over finite fields of the reduction, with respect to an equivalent one that uses a more traditional Laporta algorithm. We also break down the percentages of time spent in each main step of our method (note that they generally don’t add up to $100\%$ since additional time is needed to combine the various types of identities).
@@ Acknowledgements.
We thank Pierpaolo Mastrolia, Kay Schönwald and Simone Zoia for comments and feedback on the draft and Johann Usovitsch for discussions and suggestions on the seeding for IBP identities. This work received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme grant agreement 101019620 (ERC Advanced Grant TOPUP) and under the European Union’s Horizon Europe research and innovation programme grant agreement 101040760 (ERC Starting Grant FFHiggsTop).

### v2

@@ 1 Introduction
In this paper, we study the impact of transverse integration (TI) identities on the solution of the Laporta system of equations, and assess the performance improvement they can provide. Starting from an integral family contributing to a process, we identify integrals corresponding to diagrams with either fewer external legs than the original family, or that can be factorized into lower-loop integrals. These can be mapped to simpler integral families having fewer invariants or fewer loops, making the identities for these new families much simpler. We map these integrals to the new families via a transverse integration, namely a tensor decomposition in the subspace that is orthogonal to their external legs. While this technique already appeared, e.g. in the context of integrand decomposition Mastrolia:2016dhn and multi-loop numerical unitarity Abreu:2017xsl ; Abreu:2020xvt , its impact on a traditional IBP reduction is unexplored.
@@ 3 The main idea in a pedagogical example
Its kinematics are described by $3$ massless and one massive external momenta and we have $3$ independent external invariants, as follows
@@ 3.1 Integrals with fewer external legs
As we will see in section 4, the form factors $C_{j}$ can be obtained by contracting the tensor integral on the l.h.s. of (27) with projectors $P_{j,\perp}$ that are linear combinations of the same tensors $T_{j,\perp}$. The coefficients of this linear combinations are rational functions of $D_{\perp}$ that can be determined in a general way and only depend on the tensor rank $\beta_{1}+\beta_{2}$ and not on the integral family (see also Anastasiou:2023koq ; Goode:2024mci ). By applying these projectors to the tensor integrals, we obtain scalar products of the form $(k_{i,\perp}\cdot k_{j,\perp})$, that in turn may be rewritten as
@@ 4.1 Sectors with fewer external legs
where $C_{j}$ are scalar form factors and $T_{\perp,j}^{\mu_{1}\cdots\mu_{r}}$ is a basis of rank-$r$ covariant tensors in the $D_{\perp}$-dimensional transverse space. Such a basis can be chosen using
@@
while the others can be obtained from $T_{1,\perp}$ via independent permutations of the Lorentz indices (there are $r!/(2^{r/2}(r/2)!)$ independent tensors, without accounting for symmetries of the integrand). Here $g_{\perp}^{\mu\nu}$ is the projection of the metric tensor into the subspace that is transverse to $q_{1},\ldots,q_{D_{\parallel}}$, as defined above. The form factors $C_{j}$ are computed by contracting the l.h.s. of eq. (73) with suitable projectors $P_{j,\perp}$ such that
@@
For most phenomenologically relevant applications, the decomposition in eq. (73) is thus easy to obtain even via a naive inversion of the matrix $T$. In cases where integrals with very high-rank numerators are needed, improved tensor reduction techniques exist (see e.g. Anastasiou:2023koq ; Goode:2024mci ) and can effectively mitigate the complexity of this step.
@@ 4.2 Factorizable sectors
$I_{F;a_{1},\cdots a_{t}0\cdots 0}=\left(\int\prod_{j=1}^{\ell_{1}}d^{D}k_{j}\frac{1}{D_{1}^{a_{1}}\cdots D_{t_{1}}^{a_{t_{1}}}}\right)\left(\int\prod_{j=\ell_{1}+1}^{\ell_{1}+\ell_{2}}d^{D}k_{j}\frac{1}{D_{t_{1}+1}^{a_{t_{1}+1}}\cdots D_{t_{1}+t_{2}}^{a_{t_{1}+t_{2}}}}\right),$ (85)
@@
$I_{F;a_{1},\cdots a_{t}0\cdots 0}[N(k_{1},\ldots,k_{\ell})]=\left(\int\prod_{j=\ell_{1}+1}^{\ell_{1}+\ell_{2}}d^{D}k_{j}\frac{1}{D_{t_{1}+1}^{a_{t_{1}+1}}\cdots D_{t_{1}+t_{2}}^{a_{t_{1}+t_{2}}}}\right)I_{F_{1};a_{1}\cdots a_{t_{1}}0\cdots 0}[N(k_{1},\ldots,k_{\ell})],$ (87)
where the numerator $N$ is polynomial in the ISPs of $F$ and therefore in the scalar products one can build from the loop and external momenta of $F$. We thus focus on the $\ell_{1}$-loop integral $I_{F_{1};a_{1}\cdots a_{t_{1}}0\cdots 0}[N(k_{1},\ldots,k_{\ell})]$ and handle its numerator by following the same TI steps as in section 4.1. More specifically, $N(k_{1},\ldots,k_{\ell})$ may contain scalar products which cannot be written as a linear combination of generalized denominators of $F_{1}$. The corresponding integrals, however, will be cast – by means of the algorithm in 4.1 – as a linear combination of integrals $I_{F;\vec{b}}$ in the family $F_{1}$. Moreover, by reviewing these steps of the TI algorithm, it is easy to see that the coefficients $N_{\vec{b}}$ of such linear combination will contain, in their denominators, only scalar products between the external legs $q_{1},\ldots,q_{D_{\parallel}}$ of sector $S_{F_{1};a_{1}\cdots a_{t_{1}}0\cdots 0}$. Hence, the dependence of these coefficients on any other momentum which may appear in $N(k_{1},\ldots,k_{\ell})$, including the other loop momenta $k_{\ell_{1}+1},\ldots k_{\ell_{1}+\ell_{2}}$, will be in the numerator of these coefficients. We thus obtain a relation of the form
@@
The same method can be used, recursively, for factorizable sectors whose corner integral is the product of more than two sub-loop factors.
@@ 5.2 Implementation
These three identities will thus be combined numerically over finite fields. While this strongly reduces the size of the expressions, it doesn’t completely fix the issue for complex multi-scale integrals having high rank. This happens in particular for step 1 above, and sometimes also for step 3. We note, however, that the identities in eq. (70) and eq. (79) for individual generalized denominators or scalar products are relatively simple. Moreover, they are simply substitutions of loop variables which we can implement at the integrand level. When these are raised to high powers or multiplied to become high-rank integrals, on the other hand, expressions can still become very large. Finite-field technologies, however, allow us to fix this issue with our second trick. We first encode the expressions for the required generalized denominators and scalar products as (linear) polynomials in the loop variables appearing on the r.h.s. of the equations we use in steps 1 and 3 above. The coefficients of such polynomials, which only depend on the external variables, are instead treated numerically over finite fields. This drastically simplifies their expressions. Hence, for each numerical evaluation, we can afford to build higher powers of (products of) these polynomials via a sequence of polynomial multiplications over finite fields. At the time of writing, FiniteFlow does not expose polynomial multiplication algorithms to its Mathematica interface, so these have been implemented as a sequence of linear substitutions, for which FiniteFlow’s sparse solver was used. A more specialized and optimized algorithm might be used in the future, but this was sufficient for our applications. These two tricks thus allowed us to implement complex TI identities without ever dealing with huge expressions.
@@ 5.3 Examples
In particular, we report on the relative performance improvement of a numerical evaluation over finite fields of the reduction, with respect to an equivalent one that uses a more traditional Laporta algorithm. The latter has also been implemented using the choice of seed integrals proposed in Peraro:2019svx (also described in section 5.1) and the weight choice proposed in Appendix B of the same reference. We stress that the choice of seed integrals for our new approach and the traditional approach we compare against is thus identical, except for sectors where simplifications are enabled by the use of TI identities. Both approaches have been implemented using the Mathematica interface of the FiniteFlow program.
For all integral families we consider, we reduce the full set of integrals belonging to non-zero sectors with $u=0$ and $s\leq s_{\textrm{max}}$ for some maximum rank $s_{\textrm{max}}$ specified below. In the following, we also break down the percentages of time spent in each main step of our method (note that they generally don’t add up to $100\%$ since additional time is needed to combine the various types of identities).
@@ Acknowledgements.
We thank Pierpaolo Mastrolia, Kay Schönwald and Simone Zoia for comments and feedback on the draft and Johann Usovitsch for discussions and suggestions on the seeding for IBP identities. VC and TP received funding from the European Research Council (ERC) under the European Union’s Horizon Europe research and innovation programme grant agreement 101040760 (ERC Starting Grant FFHiggsTop), GF received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme grant agreement 101019620 (ERC Advanced Grant TOPUP) and from the UZH Candoc scheme (Candoc grant Nifty LooPS).